A brand of lightbulb has an average life of 600 hr, with a standard deviation of 50 hr. The length of life of the bulb can be closely approximated by a normal curve. A warehouse manager buys and installs​ 10,000 such bulbs. Find the total number that can be expected to last less than 730 hr.

the total number of bulbs that can be expected to last less than

730 hr is ?

​(Round to the nearest whole number as​ needed.)
Answer #1

Given,

= 600 , = 50

We convert this to standard normal as

P( X < 730) = P( Z < x - / )

= P( Z < 730 - 600 / 50 )

= P( Z < 2.6)

= 0.9953

Expected number = 10000 * 0.9953 = 9953